In​ 2012, the population of a city was 5.72 million. The exponential growth rate was 3.79 per year.
​a) Find the exponential growth function.
​b) Estimate the population of the city in 2018.
​c) When will the population of the city be 88 ​million?
​d) Find the doubling time.
This one's also a big one and it is pretty hard.

Answer by stanbon(75887)

In​ 2012, the population of a city was 5.72 million. The exponential growth rate was 3.79 per year.
​a) Find the exponential growth function.
A(t) = 5.72*3.79^t
-------------------------------
​b) Estimate the population of the city in 2018.
A(6) = 5.72*2.79^6
--------------------------------
​c) When will the population of the city be 88 ​million?
Solve for "t"::
88 = 5.72*2.79^t
2.79^t = 15.38
t = log(15.38)/(log(2.79)) = 2.66 years
Ans:: 2018 + 2.66 yrs = 2021
---------------------------
​d) Find the doubling time.
2*5.72 = 5.72*3.79^t
3.79^t = 2
t = log(2)/log(3.79)
t = 0.52
Ans:: 2018 + 0.52(12 mts) = 2018 and 6 mts
--------------
Cheers,
Stan H.
-------------
This one's also a big one and it is pretty hard.